Journal Announcement and Call for Papers for a Full-Service, Refereed Electronic Journal

The New York Journal

of Mathematics

The New York Journal of Mathematics is a refereed mathematics journal being launched by the University at Albany, State University of New York.

The journal is broadly based in subject matter, covering algebra, modern analysis, and geometry/topology.

The papers will be presented via the TEX mathematical typesetting language, developed by Donald Knuth. TEX has become the standard for computer typesetting of mathematical publications.

Access will be provided through a variety of electronic means, including listserv, anonymous ftp and gopher. In particular, access will be much more convenient than that provided by a conventional print journal. Because papers will be published as they are accepted by the editorial board, publication delays will be minimal.

The articles will be archived, date stamped, and maintained by the University at Albany, State University of New York. The University takes responsibility for insuring the integrity of the journal’s archive in perpetuity.

The journal will have an ISSN number and will be accessed by the major bibliographic publications in the field.

The development of this journal is being underwritten by the Office of Information Systems of the University at Albany, State University of New York, and is supported in part by a grant from the Office of the Vice President for Research.

The Office of Information Systems includes both the University Libraries and Computing and Network Services. These two units are migrating toward merged operations as part of the University’s Electronic Library Initiative. The journal itself is part of this initiative.

Thus, the journal is integrated with the library system, and professional librarians have been active in every phase of the journal’s development, including the provisions for the preservation of the data and the maintenance of the integrity of the archive.

Here, we list the editorial board and give details for the submission of articles and for access to the journal.

1. Editorial Board

Mark Steinberger, The University at Albany, State University of New York

Editor in Chief mark@sarah.albany.edu Marc Culler University of Illinois at Chicago culler@math.uic.edu

Andrew Granville University of Georgia nyj@sophie.math.uga.edu

Jonathan Lubin Brown University lubinj@gauss.math.brown.edu

Paul Muhly University of Iowa muhly@math.uiowa.edu

Bodo Pareigis Universitaet Muenchen^{1}

Andrew Ranicki University of Edinburgh a.ranicki@edinburgh.ac.uk Douglas C. Ravenel University of Rochester drav@troi.cc.rochester.edu

Richard Rochberg Washington University rr@math.wustl.edu

Joseph Rosenblatt Ohio State University jrsnbltt@math.ohio-state.edu

David J. Saltman University of Texas saltman@math.utexas.edu

Birgit Speh

Cornell University speh@math.cornell.edu

Ronald J. Stern

University of California at Irvine rstern@math.uci.edu

2. Instructions for Submission of Articles

Complete information for submission of articles is given in the journal’s gopher/ftp archive. We summarize that information here.

Papers should be submitted by electronic mail to the member of the editorial board closest to your field. For the convenience of authors, a guide to the areas considered by the individual editors has been placed in the journal’s archive.

The papers will be refereed in the traditional manner, with one anonymous referee.

The specifications for our articles are described explicitly in Style specifications for electronic submissions to the New York Journal of Mathematics, on file in the archive. Copies are also available from the listserv.

Two of these specifications should be emphasized: We ask that authors prepare their papers using the style files and the templates for front matter provided by either the archive or the listserv. They are designed to facilitate the editing process.

We currently have style specifications for submissions in AMS-TEX and AMS-L^{A}TEX. We may add other flavors of TEX at a later date.

We will permit the use of certain specified diagram packages and techniques for the inclusion of graphics. Details will be provided in the style specifications.

Authors will have an opportunity to make final corrections before publication.

1Pareigis@rz.Mathematik.Uni-Muenchen.de

3. Access to the New York Journal of Mathematics

The New York Journal will be accessible by a combination of listserv and gopher/ftp.

By listserv, we will distribute abstracts of the papers, together with instructions for retrieving them via ftp and gopher.

The abstracts will be available on four separate listserv lists: one will distribute abstracts for every paper. The other three are specialty lists, giving abstracts for papers in algebra, analysis and geometry/topology, respectively. The reader may, of course, subscribe to any combination of these lists.

To subscribe, send a note to listserv@albany.edu (or listserv@albany.bitnet):

If you want to receive the abstracts for all the papers, the body of your note should read

subscribe NYJMTH-A ⟨your full name⟩

For the specialized lists, the body of your note should read

Algebra:

subscribe NYJM-ALG ⟨your full name⟩

Analysis:

subscribe NYJM-AN ⟨your full name⟩

Geometry/Topology:

subscribe NYJM-TOP ⟨your full name⟩

To avoid duplication, we will only be archiving the postings on NYJMTH-A, the list carrying all the abstracts. The archive for NYJMTH-A will also contain the style specifications for submissions and the templates for frontmatter.

The listserv archive is accessible to anyone with electronic mail. One need not even be a member of the listserv list NYJMTH-A. To get a list of the files available from the listserv archive, send a message to listserv@albany.edu with the message “index NYJMTH-A”. You may then request any of the files in the index by sending a message to listserv@albany.edu with the message “get ⟨filename⟩”. Be sure to send these messages to the listserv, rather than the list itself.

The gopher/ftp archive will contain both the abstract and a .dvi file for each paper. Both abstract and .dvi will remain unchanged from the official date of publication, as may be verified by checking the date stamps.

The gopher archive may be accessed by typing “gopher nyjm.albany.edu”. It is also available under the “Academic Offerings, Departments and Programs” selection of the main gopher menu of the State University of New York at Albany.

The same archive is accessible by anonymous ftp to nyjm.albany.edu, in the directory /pub/nyjm.

The reader must have access to a printer driver for TEX in order to print out the .dvi files. Such a driver is present in most mathematics departments. Math departments, libraries, and other institutional subscribers are welcome to consult with the Editor in Chief.